OpenStudy (anonymous):

Please help me solve this :exact value of inverse tan (1/3 sq rt 3) it has been bothering me for days b/c its format is strange to me, and googling it brings no results!!!

OpenStudy (anonymous):

\[\tan^{-1} \frac{ 1 }{ 3 }\sqrt{3}\]

OpenStudy (anonymous):

"find the exact value of the following expression. If undefined stae undefined a-1/3 pi b-1/6 pi c--1/3 pi d-undefined e--1/6 pi

OpenStudy (anonymous):

*state

hartnn (hartnn):

\[\frac{1}{3}\sqrt{3}=\frac{1}{\sqrt{3}}\] \[if{\space} \tan \theta = x {\space}then {\space} \theta=\tan ^{-1}x\] so uhust need to find for what value of \theta,u get 1/root3

hartnn (hartnn):

*u just need

hartnn (hartnn):

u know how to find tan 1/3 pi tan -1/3 pi ?

OpenStudy (anonymous):

they write it so weird! -_-... i dont remember! took it 2 yrs ago..please elaborate?

hartnn (hartnn):

\[\tan ^{-1} \frac{1}{\sqrt{3}}=30 {\space} =\pi/6\] because \[\tan \pi/6 = \frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

oh..righ right, degree to radian! *face palm

hartnn (hartnn):

oh, \[\pi ^r = 180 ^0\] the conversion is 1 degree=pi/180 radians

OpenStudy (anonymous):

thank you!